OpenStudy (anonymous):

Which of the following is equivalent to three and two-fifths divided by four? one and three-seventeenths three and one-tenth fifteen twentieths seventeen-twentieths

OpenStudy (anonymous):

D, 17/5=3.4 3.4/4=0.85 17/20=0.85